OpenStudy (anonymous):

the ratio of the areas of two rectangles is 4 to 5. if the area of the larger rectangles is 125 sq feet what is the area of the smaller rectangle?

OpenStudy (anonymous):

4/5 = x/125 4*125 /5 = x 100 = x

OpenStudy (anonymous):

thanks : )